how do I solve:
g(n)= n^2 + 4 + 2n
h(n)= -3n + 2
Find (g x h)(1)

Answer by ewatrrr(24785)


Hi, Previously Posted
(g x h)(n) = -3n(n^2+2n+4) + 2(n^2+2n+4)
(g x h)(n) = -3n^3-4n^2 -8n + 8
(g x h)(1) = -3n^3-4n^2 -8n + 8 = -3 -4 -8 + 8 = -7